# SpeechAlign: a Framework for Speech Translation Alignment Evaluation

### v1

@@ Abstract
Speech-to-Speech and Speech-to-Text translation are currently dynamic areas of research. To contribute to these fields, we present SpeechAlign, a framework to evaluate the underexplored field of source-target alignment in speech models. Our framework has two core components. First, to tackle the absence of suitable evaluation datasets, we introduce the Speech Gold Alignment dataset, built upon a English-German text translation gold alignment dataset. Secondly, we introduce two novel metrics, Speech Alignment Error Rate (SAER) and Time-weighted Speech Alignment Error Rate (TW-SAER), to evaluate alignment quality in speech models. By publishing SpeechAlign we provide an accessible evaluation framework for model assessment, and we employ it to benchmark open-source Speech Translation models.
1 Introduction
Speech-to-text Translation (S2TT) and Speech-to-speech Translation (S2ST) refer to the task of converting spoken language into respectively written text or speech in a different language. This tasks are increasing their popularity, and can be used for applications such as subtitling videos in a different language, translating between languages that do not have a written form, and in general, ensuring seamless communication across people worldwide.
@@ 1 Introduction
In speech-related fields, little interpretability work regarding alignments has been done. Some previous studies have focused on analysing the self-attention in the encoder of speech recognition, Zhang et al. (2021); Shim et al. (2022) and speech translation Alastruey et al. (2022) systems. However, these models' decoder, and consequently, its alignment capabilities, have yet to be explored, potentially due to the absence of suitable datasets and metrics for evaluating the task in this setting.
In light of this, we introduce SpeechAlign framework [^1] Available on request. Link will be published upon conference acceptance., which serves as a solution to the stated lack of resources. SpeechAlign is formed of two core components: a novel dataset and an evaluation framework founded on our proposed metrics.
@@ 2 Related Work
3 Speech Gold Alignment Dataset
@@
This augmented dataset, can either be considered as two distinct datasets supporting S2TT from English to German (and vice versa), or as a unified S2ST dataset by combining both S2TT alignments. In Figure 2 we show the three different modalities of the dataset. Figure 1(a) shows a sample from the original dataset presented by Vilar et al. (2006), and Figures 1(b) and 1(c) show our extension for S2TT and S2ST settings respectively.
@@ 3.1 Methodology
3.1.1 Speech Generation
To produce synthetic speech for the sentences in the Gold Alignment dataset, we employed the VITS model. This TTS system uses a phonemizer to obtain the phonemes corresponding to the input sequence. Then, to generate the speech output, the model uses a stochastic duration predictor that assigns a duration to each phoneme. The chosen duration is randomly sampled from each phoneme's durations distribution. By doing this, the model is able to synthesize natural speech and can generate different speech utterances for the same input text.
@@ 3.1.1 Speech Generation
3.1.2 Word-Audio Matching
@@
The approach followed to accomplish this starts by acquiring intermediate representations from VITS. Specifically, we gather the output generated by the phonemizer, which is the phonemized sentence, as well as the output of the duration predictor. This predictor crates a dictionary containing duration in integer units of each phoneme. With this information in hand, we perform a two-step matching procedure, that ultimately yields the mapping from audio to words, via the intermediate representation of phonemes:
@@
3.1.3 Special Cases
During the two phases of the dataset construction we encounter some special challenges that need special handling.
@@ 3.1.3 Special Cases
Possesives ('s)
The original Gold Alignment dataset does not provide alignments between natural sentences, but for sentences tokenized with Moses. However, VITS works on natural text, and this missmatch created some difficulties along the matching process. This is the case of words such as "Parliament's", that is considered a single word when dealing with VITS ("Parliament's" $\rightarrow$ /pal@m@nts/) , but it is actually two different words with independent alignments in the original dataset ("Parliament 's"), due to Moses tokenization.
This is a case of Phonemic Fusion (and it's addressed as such). However, unlike previously shown cases caused by the phonemizer, this fusion stems from the tokenization in the original dataset.
@@
A similar behaviour arises when dealing with percent signs. These signs appear alongside numbers in natural text ("34%"), but in the Gold Alignment dataset, they're separate tokens due to Moses tokenization ("34 %"). However, as illustrated in Table 1, percents are a case of Phonemic Fragmentation, with the phonemizer breaking this construction into multiple phonemized words ("34%" $\rightarrow$ /ðEtI fO p@sent/).
In this particular cases of Phonemic Fragmentation, we aim to separate the expanded phonetic text into two segments: the first containing phonemized words associated with the number (/ðEtI fO/), and the last containing the phonemized word corresponding to the percent (/p@sent/). In this instances, the merging of time intervals encompasses all words except the final one in the expansion.
@@
In our utilization of the German phonemizer, we have noticed that it occasionally produces inaccurate phonetic transcriptions for certain single input words. These inaccuracies tend to occur with special symbols (e.g., "%", "/"), years (e.g., "1996"), acronyms (e.g., "EU", "Nr"), compound nouns (e.g., "EU-Staate"), among others. To rectify these inaccuracies in phonetic transcription, we have replaced specific words in the input sentences with their expanded and "spoken" format ("EU" $\rightarrow$ "E U", "1996" $\rightarrow$ "nineteen ninety six"). This adjustment assists the phonemizer in producing more accurate transcriptions.
3.2 Dataset Quality Assessment
@@
In Table 2, we present the Word Error Rate (WER) results obtained on both datasets, and we observe that our synthetic audios results in a lower WER than standard EuroParl ST dataset. Consequently, we can conclude that the synthesized data does not pose a problem and appears to be easily handled by the models, possibly due to the clarity of the generated audios compared to European Parliament recordings.
@@
4 Proposed Evaluation
The objective of this section is to define an evaluation procedure and metrics that are able assess models' ability to establish source-target alignments. To analyse this capability, our focus is on the contribution maps generated by the models. These maps indicate the relationship between source and target tokens, such that the contribution of a source token to a target one is always a non-negative value, and that the sum of contributions from all source tokens to a target token must equal 1 (i.e. attention weights or more advanced interpretability methods Kobayashi et al. (2021); Ferrando et al. (2022)). Then, to measure the alignments, we build new metrics around the intuition of the Alignment Error Rate (AER) score, initially introduced by Och and Ney (2003) and defined in section 2. However, extracting the alignments from the contribution map and adapting AER for speech sequences is not straightforward process.
@@ 4 Proposed Evaluation
Input: $C\_t2t$: token-to-token contribution matrix,
@@
$max\_duration\_src\leftarrow src[-1][end]$ $max\_tokens\_src\leftarrow C\_t2t.shape[1]$ for $word,word\_idx\leftarrow src$ do
    $s\_time\leftarrow src[word][start]$ $e\_time\leftarrow src[word][end]$ $s\_token\leftarrow ceil(s\_time*max\_tokens\_src/max\_duration\_src)$ $e\_token\leftarrow floor(e\_time*max\_tokens\_src/max\_duration\_src)$ $C\_w2t[:,word\_idx]\leftarrow\text{sum}(C\_t2t[:,s\_token:e\_token],dim=1)$
    $max\_duration\_tgt\leftarrow tgt[-1][end]$ $max\_tokens\_tgt\leftarrow C\_t2t.shape[0]$ for $word,word\_idx\leftarrow tgt$ do
       $s\_time\leftarrow tgt[word][start]$ $e\_time\leftarrow tgt[word][end]$ $s\_token\leftarrow ceil(s\_time*max\_tokens\_tgt/max\_duration\_tgt)$ $e\_token\leftarrow floor(e\_time*max\_tokens\_tgt/max\_duration\_tgt)$ $C\_w2w[word\_idx,:]\leftarrow\text{avg}(C\_w2t[s\_token:e\_token,:],dim=0)$
@@ 4.1 Preprocessing
4.2 Speech Alignment Error Rate
Once we have the hard alignments, we define the Speech Alignment Error Rate (SAER) in the same manner the AER is defined. This is, given a set set $S$ of unambiguous alignments, a set $P$ of ambiguous alignment and a set $A$ of hypothesis alignment:
$SAER=1-\frac{|A\cap S|+|A\cap P|}{|A|+|S|}$ (2)
However, it's important to note that SAER doesn't fully address a key aspect in the speech setting – the noticeable disparity in the number of different tokens that form each word, which corresponds to audio durations. Instead, when computing SAER each word contributes equally to the final score, regardless of its duration.
4.3 Time-Weighted SAER
$TW-SAER=1-\frac{\sum_{i,j\in A\cap S}w_{i,j}+\sum_{i,j\in A\cap P}w_{i,j}}{\sum_{i,j\in A}w_{i,j}+\sum_{i,j\in S}w_{i,j}}$ (4)
Figure 6: TW-SAER weights.
5 SpeechAlign
@@
Finally, using SpeechAlign, we benchmark some S2TT models. For simplicity, we decide to analyze alignments based on models' cross-attention weights. We decide not benchmark the S2ST task due to the current lack of open-source models, being the recently published SeamlessM4T SeamlessCommunication et al. (2023) the only one available as of now. This model comprises two consecutive Transformers, each containing its own decoder. Consequently, it presents significant challenges in terms of obtaining a contributions map based on attention weights, and developing further interpretability methods lies beyond scope of this paper.
@@
Table 3 presents an evaluation of various sizes of the Whisper model Radford et al. (2022) on De-En S2TT. Each model's performance is assessed through the BLEU score on our test set, and the SAER and TW-SAER. The latter are computed on the attention weights of each decoder layer, and in Table 3 we report the best obtained score. This analysis uncovers a correlation between the performance metrics and the alignment score. This correlation is also observed to align with the model's size. Intriguingly, an outlier appears in the case of the Large V2 model, which fails to deliver proper alignment despite achieving the highest performance metrics.
6 Conclusion
@@
In conclusion, this paper introduces SpeechAlign, a framework to evaluate alignment in speech models. SpeechAlign has two main components. Firstly, we've created the Speech Gold Alignment dataset, being the first of its kind and created to address the lack of suitable evaluation data for the task. Secondly, we have presented the two first evaluation metrics for speech alignment, Speech Alignment Error Rate (SAER) and Time-weighted Speech Alignment Error Rate (TW-SAER), to assess how well speech models perform on the alignment task. SpeechAlign provides an accessible way to evaluate speech models, and we have used it to benchmark various open-source models.
Limitations
Our study only focuses on high resource languages, German and English, stemming from the limitation of the original Gold Alignment dataset Vilar et al. (2006). Additionally, it's worth noting the potential challenges that might arise due to the limited domain covered by the dataset, centered on speeches from the European Parliament. This might cause some issues when evaluating models trained with a limited amount of data and from other domains. However, this case would be a limitation of the model itself, denoting its restricted capabilities. Using data from European Parliament's speeches is a common practice in the Machine Translation field Koehn (2005); Iranzo-Sánchez et al. (2020).

### v2

@@ Abstract
Speech-to-Speech and Speech-to-Text translation are currently dynamic areas of research. In our commitment to advance these fields, we present SpeechAlign, a framework designed to evaluate the underexplored field of source-target alignment in speech models. The SpeechAlign framework has two core components. First, to tackle the absence of suitable evaluation datasets, we introduce the Speech Gold Alignment dataset, built upon a English-German text translation gold alignment dataset. Secondly, we introduce two novel metrics, Speech Alignment Error Rate (SAER) and Time-weighted Speech Alignment Error Rate (TW-SAER), which enable the evaluation of alignment quality within speech models. While the former gives equal importance to each word, the latter assigns weights based on the length of the words in the speech signal. By publishing SpeechAlign we provide an accessible evaluation framework for model assessment, and we employ it to benchmark open-source Speech Translation models. In doing so, we contribute to the ongoing research progress within the fields of Speech-to-Speech and Speech-to-Text translation.
Keywords: Evaluation Methodologies,Speech Resource/Database, SpeechToSpeech Translation
[…]
SpeechAlign: a Framework for Speech Translation Alignment Evaluation
Belen Alastruey1∗thanks:   Equal contribution. , Aleix Sant2∗, Gerard I. Gállego2, David Dale1, Marta R. Costa-jussà1
Meta FAIR, Paris, France1
TALP Research Center, Universitat Politècnica de Catalunya, Barcelona, Spain2
alastruey@meta.com, aleix.sant@estudiantat.upc.edu,
gerard.i.gallego@upc.edu, daviddale@meta.com, costajussa@meta.com
Abstract content
1.   Introduction
Speech-to-text Translation (S2TT) and Speech-to-speech Translation (S2ST) refer to the task of converting spoken language into respectively written text or speech in a different language. These tasks are increasing their popularity, and can be used for applications such as subtitling videos in a different language, translating between languages that do not have a written form, and in general, ensuring seamless communication across people worldwide.
@@ 1.   Introduction
In speech-related fields, little interpretability work regarding alignments has been done. Some previous studies have focused on analysing the self-attention in the encoder of speech recognition, Zhang et al. (2021); Shim et al. (2022) and speech translation Alastruey et al. (2022) systems. However, these models’ decoder, and consequently, its alignment capabilities, have yet to be explored, potentially due to the absence of suitable datasets and metrics for evaluating the task in this setting.
In light of this, we introduce SpeechAlign framework [^1] https://github.com/mt-upc/speechalign, which serves as a solution to the stated lack of resources. SpeechAlign is formed of two core components: a novel dataset and an evaluation framework founded on our proposed metrics.
@@ 2.   Related Work
3.   Speech Gold Alignment Dataset
@@
This augmented dataset, can either be considered as two distinct datasets supporting S2TT from English to German (and vice versa), or as a unified S2ST dataset by combining both S2TT alignments. In Figure 2 we show the three different modalities of the dataset. Figure 2(a) shows a sample from the original dataset presented by Vilar et al. (2006), and Figures 2(b) and 2(c) show our extension for S2TT and S2ST settings respectively.
@@ 3.1.   Methodology
3.1.1.   Speech Generation
To produce synthetic speech for the sentences in the Gold Alignment dataset, we employed the VITS model. This TTS system uses a phonemizer to obtain the phonemes corresponding to the input sequence. Then, to generate the speech output, the model uses a stochastic duration predictor that assigns a duration to each phoneme. The chosen duration is randomly sampled from each phoneme’s durations distribution. By doing this, the model is able to synthesize natural speech and can generate different speech utterances for the same input text.
@@ 3.1.1.   Speech Generation
3.1.2.   Word-Audio Matching
@@
The approach followed to accomplish this starts by acquiring intermediate representations from VITS. Specifically, we gather the output generated by the phonemizer, which is the phonemized sentence, as well as the output of the duration predictor. This predictor creates a dictionary containing duration in integer units of each phoneme. With this information in hand, we perform a two-step matching procedure, that ultimately yields the mapping from audio to words, via the intermediate representation of phonemes:
@@
3.1.3.   Special Cases
After the two phases of the dataset construction we perform a manual revision of the generated data and encounter some special challenges that need special handling.
@@ 3.1.3.   Special Cases
Possesives (’s)
[…]
This is a case of Phonemic Fusion (and it’s addressed as such). However, unlike previously shown cases caused by the phonemizer, this fusion stems from the tokenization in the original dataset.
@@
[…]
@@
In our utilization of the German phonemizer, we have noticed that it occasionally produces inaccurate phonetic transcriptions for certain single input words. These inaccuracies tend to occur with special symbols (e.g., ”%”, ”/”), years (e.g., ”1996”), acronyms (e.g., ”EU”, ”Nr”), compound nouns (e.g., ”EU-Staate”), among others. To rectify these inaccuracies in phonetic transcription, we have replaced specific words in the input sentences with their expanded and ”spoken” format (”EU” $\rightarrow$ ”E U”, ”1996” $\rightarrow$ ”nineteen ninety six”). This adjustment assists the phonemizer in producing more accurate transcriptions.
3.2.   Dataset Quality Assessment
@@
In Table 2, we present the Word Error Rate (WER) results obtained on both datasets, and we observe that our synthetic audios result in a lower WER than standard EuroParl ST dataset. Furthermore, we notice a disparity in performance between the German and English synthetic data. This discrepancy may stem from differences in the underlying VITS models, that are trained on distinct datasets for each language. This quality discrepancy between the German and English outputs was confirmed during the manual inspection of the generated speech. Despite this variance, it’s important to point out that WER for German remains below the threshold established by Europarl ST, which serves as a quality reference in our study. Consequently, we can conclude that the synthesized data does not pose a problem and appears to be easily handled by the models in both languages, possibly due to the clarity of the generated audios compared to European Parliament recordings.
@@
4.   Proposed Evaluation
The objective of this section is to define an evaluation procedure and metrics that are able assess models’ ability to establish source-target alignments. To analyse this capability, our focus is on the contribution maps generated by the models. These maps indicate the relationship between source and target tokens, such that the contribution of a source token to a target one is always a non-negative value, and that the sum of contributions from all source tokens to a target token must equal 1 (i.e. attention weights or more advanced interpretability methods Kobayashi et al. (2021); Ferrando et al. (2022)). Then, to measure the alignments, we build new metrics around the intuition of the Alignment Error Rate (AER) score, initially introduced by Och and Ney (2003) and defined in section 2. However, extracting the alignments from the contribution map and adapting AER for speech sequences is not straightforward process.
@@ 4.   Proposed Evaluation
Input:
$C\_t2t$: token-to-token contribution matrix,
@@
$max\_duration\_src\leftarrow src[-1][end]$
$max\_tokens\_src\leftarrow C\_t2t.shape[1]$
for $word,word\_idx\leftarrow src$ do
    $s\_time\leftarrow src[word][start]$
    $e\_time\leftarrow src[word][end]$
    $s\_token\leftarrow ceil(s\_time*max\_tokens\_src/max\_duration\_src)$
    $e\_token\leftarrow floor(e\_time*max\_tokens\_src/max\_duration\_src)$
    $C\_w2t[:,word\_idx]\leftarrow\text{sum}(C\_t2t[:,s\_token:e\_token],dim=1)$
    $max\_duration\_tgt\leftarrow tgt[-1][end]$
    $max\_tokens\_tgt\leftarrow C\_t2t.shape[0]$
    for $word,word\_idx\leftarrow tgt$ do
       $s\_time\leftarrow tgt[word][start]$
       $e\_time\leftarrow tgt[word][end]$
       $s\_token\leftarrow ceil(s\_time*max\_tokens\_tgt/max\_duration\_tgt)$
       $e\_token\leftarrow floor(e\_time*max\_tokens\_tgt/max\_duration\_tgt)$
       $C\_w2w[word\_idx,:]\leftarrow\text{avg}(C\_w2t[s\_token:e\_token,:],dim=0)$
@@ 4.1.   Preprocessing
4.2.   Speech Alignment Error Rate
Once we have the hard alignments, we define the Speech Alignment Error Rate (SAER) in the same manner the AER is defined. This is, given a set $S$ of unambiguous alignments, a set $P$ of ambiguous alignment and a set $A$ of hypothesis alignment:
$SAER=1-\frac{|A\cap S|+|A\cap P|}{|A|+|S|}$ (2)
Note that while the equation remains identical to that in the original AER case, the definitions of $A$, $S$, and $P$ change due to the preprocessing required to derive them.
Furthermore, it’s important to note that SAER doesn’t fully address a key aspect in the speech setting – the noticeable disparity in the number of different tokens that form each word, which corresponds to audio durations. Instead, when computing SAER each word contributes equally to the final score, regardless of its duration. Hence, tokens that correspond to short words are assigned a higher weight in the metric than those that correspond to longer words. This differs from the model perspective, where each token holds equal importance.
4.3.   Time-Weighted SAER
$\small TW-SAER=1-\frac{\sum_{i,j\in A\cap S}w_{i,j}+\sum_{i,j\in A\cap P}w_{i,j}}{\sum_{i,j\in A}w_{i,j}+\sum_{i,j\in S}w_{i,j}}$ (4)
Figure 6: TW-SAER weights.
@@ 4.3.   Time-Weighted SAER
5.   SpeechAlign
@@
Finally, using SpeechAlign, we benchmark some S2TT models. For simplicity, we decide to analyze alignments based on models’ cross-attention weights. We decide not benchmark the S2ST task due to the current lack of open-source models, being the recently published SeamlessM4T SeamlessCommunication et al. (2023) the only one available as of now. This model comprises two consecutive Transformers, each containing its own decoder. Consequently, it presents significant challenges in terms of obtaining a contributions map based on attention weights, and developing further interpretability methods lies beyond scope of this paper.
@@
Table 3 presents an evaluation of various sizes of the Whisper model Radford et al. (2022) on De-En S2TT. Each model’s performance is assessed through the BLEU score on our test set, and the SAER and TW-SAER. The latter are computed on the attention weights of each decoder layer, and in Table 3 we report the best obtained score. This analysis uncovers a correlation between the performance metrics and the alignment score. This correlation is also observed to align with the model’s size.
6.   Conclusion
@@
In conclusion, this paper introduces SpeechAlign, a framework to evaluate alignment in speech models. SpeechAlign has two main components. Firstly, we’ve created the Speech Gold Alignment dataset, being the first of its kind and created to address the lack of suitable evaluation data for the task. Secondly, we have presented the two first evaluation metrics for speech alignment, Speech Alignment Error Rate (SAER) and Time-weighted Speech Alignment Error Rate (TW-SAER), to assess how well speech models perform on the alignment task. SpeechAlign provides an accessible way to evaluate speech models, and we have used it to benchmark various open-source models.
7.   Bibliographical References
[…]
  • Agarwal et al. (2023) Milind Agarwal, Sweta Agrawal, Antonios Anastasopoulos, Luisa Bentivogli, Ondřej Bojar, Claudia Borg, Marine Carpuat, Roldano Cattoni, Mauro Cettolo, Mingda Chen, William Chen, Khalid Choukri, Alexandra Chronopoulou, Anna Currey, Thierry Declerck, Qianqian Dong, Kevin Duh, Yannick Estève, Marcello Federico, Souhir Gahbiche, Barry Haddow, Benjamin Hsu, Phu Mon Htut, Hirofumi Inaguma, Dávid Javorský, John Judge, Yasumasa Kano, Tom Ko, Rishu Kumar, Pengwei Li, Xutai Ma, Prashant Mathur, Evgeny Matusov, Paul McNamee, John P. McCrae, Kenton Murray, Maria Nadejde, Satoshi Nakamura, Matteo Negri, Ha Nguyen, Jan Niehues, Xing Niu, Atul Kr. Ojha, John E. Ortega, Proyag Pal, Juan Pino, Lonneke van der Plas, Peter Polák, Elijah Rippeth, Elizabeth Salesky, Jiatong Shi, Matthias Sperber, Sebastian Stüker, Katsuhito Sudoh, Yun Tang, Brian Thompson, Kevin Tran, Marco Turchi, Alex Waibel, Mingxuan Wang, Shinji Watanabe, and Rodolfo Zevallos. 2023. FINDINGS OF THE IWSLT 2023 EVALUATION CAMPAIGN. In Proceedings of the 20th International Conference on Spoken Language Translation (IWSLT 2023), pages 1–61, Toronto, Canada (in-person and online). Association for Computational Linguistics.
  • Alastruey et al. (2022) Belen Alastruey, Javier Ferrando, Gerard I. Gállego, and Marta R. Costa-jussà. 2022. On the locality of attention in direct speech translation. In Proceedings of the 60th Annual Meeting of the Association for Computational Linguistics: Student Research Workshop, pages 402–412, Dublin, Ireland. Association for Computational Linguistics.
  • Ansari et al. (2020) Ebrahim Ansari, Amittai Axelrod, Nguyen Bach, Ondrej Bojar, Roldano Cattoni, Fahim Dalvi, Nadir Durrani, Marcello Federico, Christian Federmann, Jiatao Gu, Fei Huang, Kevin Knight, Xutai Ma, Ajay Nagesh, Matteo Negri, Jan Niehues, Juan Pino, Elizabeth Salesky, Xing Shi, Sebastian Stuker, Marco Turchi, Alex Waibel, and Changhan Wang. 2020. Findings of the iwslt 2020 evaluation campaign. In IWSLT 2020.
  • Bentivogli et al. (2021) Luisa Bentivogli, Mauro Cettolo, Marco Gaido, Alina Karakanta, Alberto Martinelli, Matteo Negri, and Marco Turchi. 2021. Cascade versus direct speech translation: Do the differences still make a difference? CoRR, abs/2106.01045.
  • Eren and The Coqui TTS Team (2021) Gölge Eren and The Coqui TTS Team. 2021. Coqui TTS.
  • Ferrando et al. (2022) Javier Ferrando, Gerard I. Gállego, Belen Alastruey, Carlos Escolano, and Marta R. Costa-jussà. 2022. Towards opening the black box of neural machine translation: Source and target interpretations of the transformer. In Proceedings of the 2022 Conference on Empirical Methods in Natural Language Processing, pages 8756–8769, Abu Dhabi, United Arab Emirates. Association for Computational Linguistics.
  • Garg et al. (2019) Sarthak Garg, Stephan Peitz, Udhyakumar Nallasamy, and Matthias Paulik. 2019. Jointly learning to align and translate with transformer models. In Proceedings of the 2019 Conference on Empirical Methods in Natural Language Processing and the 9th International Joint Conference on Natural Language Processing (EMNLP-IJCNLP), pages 4453–4462, Hong Kong, China. Association for Computational Linguistics.
  • Ghader and Monz (2017) Hamidreza Ghader and Christof Monz. 2017. What does attention in neural machine translation pay attention to? In Proceedings of the Eighth International Joint Conference on Natural Language Processing (Volume 1: Long Papers), pages 30–39, Taipei, Taiwan. Asian Federation of Natural Language Processing.
  • Graça et al. (2008) João Graça, Joana Paulo Pardal, Luísa Coheur, and Diamantino Caseiro. 2008. Building a golden collection of parallel multi-language word alignment.
  • Holmqvist and Ahrenberg (2011) Maria Holmqvist and Lars Ahrenberg. 2011. A gold standard for English-Swedish word alignment.
  • Iranzo-Sánchez et al. (2020) Iranzo-Sánchez, Javier and Silvestre-Cerdà, Joan Albert and Jorge, Javier and Roselló, Nahuel and Giménez, Adrià and Sanchis, Albert and Civera, Jorge and Juan, Alfons. 2020. Europarl-ST: A Multilingual Corpus for Speech Translation of Parliamentary Debates.
  • Ito and Johnson (2017) Ito, Keith and Johnson, Linda. 2017. The LJ Speech Dataset.
  • Kalchbrenner et al. (2018) Nal Kalchbrenner, Erich Elsen, Karen Simonyan, Seb Noury, Norman Casagrande, Edward Lockhart, Florian Stimberg, Aäron van den Oord, Sander Dieleman, and Koray Kavukcuoglu. 2018. Efficient neural audio synthesis. CoRR, abs/1802.08435.
  • Kim et al. (2021) Jaehyeon Kim, Jungil Kong, and Juhee Son. 2021. Conditional variational autoencoder with adversarial learning for end-to-end text-to-speech.
  • Kobayashi et al. (2020) Goro Kobayashi, Tatsuki Kuribayashi, Sho Yokoi, and Kentaro Inui. 2020. Attention is not only a weight: Analyzing transformers with vector norms. In Proceedings of the 2020 Conference on Empirical Methods in Natural Language Processing (EMNLP), pages 7057–7075. Association for Computational Linguistics.
  • Kobayashi et al. (2021) Goro Kobayashi, Tatsuki Kuribayashi, Sho Yokoi, and Kentaro Inui. 2021. Incorporating Residual and Normalization Layers into Analysis of Masked Language Models. In Proceedings of the 2021 Conference on Empirical Methods in Natural Language Processing, pages 4547–4568, Online and Punta Cana, Dominican Republic. Association for Computational Linguistics.
  • Koehn (2005) Koehn, Philipp. 2005. Europarl: A Parallel Corpus for Statistical Machine Translation.
  • Kruijff-Korbayová et al. (2006) Ivana Kruijff-Korbayová, Klára Chvátalová, and Oana Postolache. 2006. Annotation guidelines for Czech-English word alignment. In Proceedings of the Fifth International Conference on Language Resources and Evaluation (LREC’06), Genoa, Italy. European Language Resources Association (ELRA).
  • Lambert et al. (2005) Patrik Lambert, Adrià De Gispert, Rafael Banchs, and José B. Mariño. 2005. Guidelines for word alignment evaluation and manual alignment. Language Resources and Evaluation, 39(4):267–285.
  • Macken (2010) Lieve Macken. 2010. An annotation scheme and gold standard for Dutch-English word alignment. In Proceedings of the Seventh International Conference on Language Resources and Evaluation (LREC’10), Valletta, Malta. European Language Resources Association (ELRA).
  • Müller and Kreutz (2021) Müller, Thorsten and Kreutz, Dominik. 2021. Thorsten - Open German Voice (Neutral) Dataset. Zenodo. Please use it to make the world a better place for whole humankind.
  • Ney (1999) H. Ney. 1999. Speech translation: coupling of recognition and translation. In 1999 IEEE International Conference on Acoustics, Speech, and Signal Processing. Proceedings. ICASSP99 (Cat. No.99CH36258), volume 1, pages 517–520 vol.1.
  • Och and Ney (2003) Franz Josef Och and Hermann Ney. 2003. A systematic comparison of various statistical alignment models. Computational Linguistics, 29(1):19–51.
  • Radford et al. (2022) Alec Radford, Jong Wook Kim, Tao Xu, Greg Brockman, Christine McLeavey, and Ilya Sutskever. 2022. Robust speech recognition via large-scale weak supervision.
  • SeamlessCommunication et al. (2023) SeamlessCommunication, Loïc Barrault, Yu-An Chung, Mariano Cora Meglioli, David Dale, Ning Dong, Paul-Ambroise Duquenne, Hady Elsahar, Hongyu Gong, Kevin Heffernan, John Hoffman, Christopher Klaiber, Pengwei Li, Daniel Licht, Jean Maillard, Alice Rakotoarison, Kaushik Ram Sadagopan, Guillaume Wenzek, Ethan Ye, Bapi Akula, Peng-Jen Chen, Naji El Hachem, Brian Ellis, Gabriel Mejia Gonzalez, Justin Haaheim, Prangthip Hansanti, Russ Howes, Bernie Huang, Min-Jae Hwang, Hirofumi Inaguma, Somya Jain, Elahe Kalbassi, Amanda Kallet, Ilia Kulikov, Janice Lam, Daniel Li, Xutai Ma, Ruslan Mavlyutov, Benjamin Peloquin, Mohamed Ramadan, Abinesh Ramakrishnan, Anna Sun, Kevin Tran, Tuan Tran, Igor Tufanov, Vish Vogeti, Carleigh Wood, Yilin Yang, Bokai Yu, Pierre Andrews, Can Balioglu, Marta R. Costa-jussà, Onur Celebi, Maha Elbayad, Cynthia Gao, Francisco Guzmán, Justine Kao, Ann Lee, Alexandre Mourachko, Juan Pino, Sravya Popuri, Christophe Ropers, Safiyyah Saleem, Holger Schwenk, Paden Tomasello, Changhan Wang, Jeff Wang, and Skyler Wang. 2023. Seamlessm4t—massively multilingual & multimodal machine translation. arXiv.
  • Shen et al. (2017) Jonathan Shen, Ruoming Pang, Ron J. Weiss, Mike Schuster, Navdeep Jaitly, Zongheng Yang, Zhifeng Chen, Yu Zhang, Yuxuan Wang, R. J. Skerry-Ryan, Rif A. Saurous, Yannis Agiomyrgiannakis, and Yonghui Wu. 2017. Natural TTS synthesis by conditioning wavenet on mel spectrogram predictions. CoRR, abs/1712.05884.
  • Shim et al. (2022) Kyuhong Shim, Jungwook Choi, and Wonyong Sung. 2022. Understanding the role of self attention for efficient speech recognition. In International Conference on Learning Representations.
  • van den Oord et al. (2016) Aäron van den Oord, Sander Dieleman, Heiga Zen, Karen Simonyan, Oriol Vinyals, Alex Graves, Nal Kalchbrenner, Andrew W. Senior, and Koray Kavukcuoglu. 2016. Wavenet: A generative model for raw audio. CoRR, abs/1609.03499.
  • Veaux et al. (2013) Veaux, Christophe and Yamagishi, Junichi and King, Simon. 2013. The voice bank corpus: Design, collection and data analysis of a large regional accent speech database.
  • Vilar et al. (2006) David Vilar, Maja Popović, and Hermann Ney. 2006. AER: do we need to “improve” our alignments? In Proc. International Workshop on Spoken Language Translation (IWSLT 2006), pages 205–212.
  • Yamagishi et al. (2019) Yamagishi, Junichi and Veaux, Christophe and MacDonald, Kirsten. 2019. CSTR VCTK Corpus: English Multi-speaker Corpus for CSTR Voice Cloning Toolkit (version 0.92). University of Edinburgh. The Centre for Speech Technology Research (CSTR).
  • Zhang et al. (2021) Shucong Zhang, Erfan Loweimi, Peter Bell, and Steve Renals. 2021. On the usefulness of self-attention for automatic speech recognition with transformers. In 2021 IEEE Spoken Language Technology Workshop (SLT), pages 89–96.